Two stones are projected from the top of a tower in opposite directions with the same velocity V at angles 30°,60° to the horizontal. What is the relative velocity of one stone with respect to other stone?
1
Expert's answer
2019-08-05T12:06:24-0400

Considering stone 1 going in positive "x-axis" and stone 2 going in negative direction of "x-axis"

The vertical velocity of both stones is in same direction(+"ve" ).

Stone 1:

Angle with horizontal="30\degree"

Horizontal velocity("h_1" )="Vcos\theta=Vcos 30\degree
=+\frac{\sqrt{3}}{2}V \hat{i}"

Vertical velocity("v_1" )"=Vsin\theta=Vsin 30\degree=+\frac{1}{2}V\hat{j}"

Net velocity of stone 1="V_1=+\frac{\sqrt{3}}{2}V\hat{i}+\frac{1}{2}\hat{j}"

Stone 2:

Angle with horizontal="60\degree"

Horizontal velocity="Vcos\theta=Vcos 60\degree=-\frac{1}{2}V\hat{i}"

Vertical velocity"=Vsin\u03b8=Vsin60\u00b0=+\\frac{\\sqrt{3}}{2}V\\hat{j}\n\u200b"

Net velocity of stone 2="V_2=-\frac{V}{2}\hat{i}+\frac{\sqrt{3}}{2}\hat{j}"

Relative velocity ="V_{12}=V_1-V_2"

="(\frac{\sqrt{3}}{2}V+\frac{V}{2})\hat{i}+(\frac{V}{2}-\frac{\sqrt{3}V}{2})\hat{j}"